# The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release

### v1

@@ 3.2 Observational Challenges and a Method to Reach
There are important reasons why pushing for F200W () dropouts is substantially harder than F115W ( or F150W () dropouts: 1) Such galaxies are fainter and rarer. 2) The frequency of extreme emission-line galaxies increases with redshift, and JWST clearly has shown a substantial population of them at , where they can create SEDs that differ from those of purely stellar populations, boosting the rest-optical and enhancing the apparent Balmer drop. 3) The dichroic gap in NIRCam leaves a wider spacing between F200W and F277W that makes it harder to exclude a mildly reddened Balmer-drop galaxy at , particularly with [OIII] potentially enhancing F277W. 4) Many Cycle 1 programs include only F277W, F356W, F410M, and F444W.111Notable exceptions are JEMS (Williams et al., 2023) and CANUCS (Withers et al., 2023). However, the filters beyond 4 microns have reduced sensitivity due to increased zodiacal foreground, so faint galaxies may only be well-detected in the bluer two bands, which in turn can be filled with [OIII] and H. These higher-redshift break candidates are less compelling than lower redshift examples that have 3 or 4 sensitive bands redward of the proposed break.
@@
Whereas Cycle 1 programs pushed deep in F115W and/or F150W, in program 3215 we adopt a strategy optimized to identify candidates at . In particular, this search must 1) push very deep in the 2 micron region; 2) increase the spectral resolution to isolate the dropout; and 3) include additional filters in the 2.5–4 micron region to eliminate contamination from mid-redshift emission-line galaxies. We believe that this combination of high-SNR detections, well-sampled dropout features, and systematic control over contaminants can reach the high standard of proof that claims of galaxies will require. The new imaging will achieve these requirements by combining the usual NIRCam wide-band filters with well-chosen medium-band filters. By observing very deep in F182M, F210M, and F250M, the photometry can isolate a sharp break within a narrower and hence more distinctive region. Further, using the combination of medium and wide-bands in the long-wave (LW) channels will reveal strong emission lines that could supply false positive Balmer breaks.
@@
We illustrate this strategy by considering the candidate in the CEERS field (Donnan et al., 2023), scaled down by a factor of 30 in flux and assuming the noise levels of programs 1210 & 3215. While now proved to be at (Arrabal Haro et al., 2023a), this object was a plausible candidate, showing a drop in F200W by a factor of 4 relative to F277W, no detection in F150W, and a blue continuum in the LW filters. However, this photometry was also a match to models at with strong optical emission lines and reddening suppression in the UV (Naidu et al., 2022a). Figure 2 shows how these two hypotheses are starkly differentiated by the NIRCam medium bands: the Lyman break is much steeper than a dust-enhanced Balmer break, and the medium bands easily distinguish continuum from emission lines. Shifting the true galaxy in redshift and repeating the fits, Figure 3 demonstrates that this suite of filters provides very effective recovery of candidates at between 13 and 20.
@@ 3.3 Filters, Exposure Times, and Depths
The driving requirement is to go deep in F182M and F210M, as potential dropout bands. In total, F182M, F200W, and F210M will reach 91 hrs of exposure time, far more than the 11 hrs that 1210 & 1180 provide and much deeper than any other Cycle 1 program (to our knowledge, program 2079 being deepest with 17 hrs in F200W). We split the time roughly evenly between these two medium filters, as we do not know at what redshift a candidate might appear.
@@
By placing these observations on the parallel field created by program 1210, these new data build on the 55 hours of imaging depth already invested here. Given the steep number counts of galaxies at the redshift frontier, it is optimal to maximize the detection rate by integrating deeply on a single NIRCam pointing. The region in and around the HUDF in GOODS-S has the deepest and most comprehensive imaging and spectroscopy on the sky. While galaxies could be found anywhere, we site this program in GOODS-S to utilize and build upon the amazing legacy value of this field, while utilizing a position angle that allows ultra-deep MSA spectroscopy on the HUDF at a time of low zodiacal background.
@@ 3.4 Coordinated Parallels with NIRSpec
| GOODS-S Year 2 | 3215 | 1 | 5 | 228 | 168 | 42 | 168 | ||
Program 3215 follows the path of 1210, using the NIRSpec MSA as the prime instrument and NIRCam as the coordinated parallel. This results in exceptionally deep spectroscopy on the HUDF and JADES Deep mosaic, about 104 hrs of total exposure, following on the 55 hrs of program 1210 spectroscopy released in Bunker et al. (2023a).
Following the program goal of exploring the high-redshift universe, we optimize this time toward the study of high-redshift galaxies. Unlike program 1210, which was largely targeted from HST ACS optical and WFC3 infrared imaging (Bunker et al., 2023a) save for four early high-redshift candidates from JADES (Robertson et al., 2023; Curtis-Lake et al., 2023), program 3215 will largely utilize NIRCam-selected targets. This allows a great expansion in the number of targets, as the JADES data is comfortably deeper than even the deepest WFC3 HUDF data and much deeper than the flanking CANDELS imaging.
Bringing ultra-deep spectroscopy to these faint high-redshift galaxies addresses many windows to understanding the physical processes associated with the early galaxy formation. We split the spectroscopic time across three dispersing modes—up to 168 ks on source with G395M (far deeper than program 1210), 168 ks with the Prism, and 42 ks with G140M—to serve a variety of key science drivers in studies of these galaxies.
@@
Finally, we seek to characterize the role of neutral hydrogen in and around the galaxies in the reionization epoch (e.g. Jones et al., 2023a; Saxena et al., 2023; Witstok et al., 2023) by measuring the Ly emission line. The PRISM resolution is particularly coarse at the blue end of the spectrum, such that the grating substantially outperforms it for the detection of narrow lines. Further, because of the low resolution, the prism does not distinguish well between mild Ly emission and the Ly damping wing. The medium-deep G140M/F070LP spectra are therefore required and adequate to measure the Ly narrow line better than the prism and protect the damping wing analyses.
The ultradeep spectra delivered by this program will be uniquely suited to pursue several key science cases, as discussed in the following.
The G395M data will map the early mass-metallicity relation (MZR) at , reaching the unexplored low-mass regime at . Indeed, although major progress has been made with JWST in measuring the MZR and the Fundamental Metallicity Relation at z5 (e.g. Curti et al., 2023b, a; Nakajima et al., 2023), the low mass regime remains still a largely uncharted territory, except for rare, highly lensed cases (Vanzella et al., 2023) or cases where masses are not yet robustly determined (Maseda et al., 2023). With the ultra-deep spectra, we expect to accurately measure the metallicity via the detection of auroral lines, such as [OIII]4363 (Laseter et al., 2023; Sanders et al., 2023a) for tens of galaxies, and via strong line diagnostics (re-calibrated with the auroral-detected spectra) for fainter galaxies, possibly revealing signatures of Population III (Nakajima & Maiolino, 2022; Wang et al., 2022; Maiolino et al., 2023b). The spectra are also expected to provide good detections of the CIII]1907,1909 and OIII]1661,1666 transitions, which will be used to determine the C/O abundance ratio, which (due to the different enrichment timescales of carbon and oxygen) is an important clock for the galaxy star formation history (e.g. Maiolino & Mannucci, 2019; Arellano-Córdova et al., 2022; Jones et al., 2023b).
Recent results, based on a few tens of galaxies, have shown that the gas excitation increases steeply at . This trend is likely a consequence of higher ionization parameter, lower metallicity, and a harder radiation field, possibly resulting from stellar -enhancement or contribution from X-ray binaries, or even an excess of cosmic rays (Trump et al., 2023; Cameron et al., 2023; Fujimoto et al., 2023b; Katz et al., 2023; Sanders et al., 2023b; Trussler et al., 2022). The 3215 deep spectra will explore these trends towards lower mass galaxies, constraining the physics of the ISM with unprecedented accuracy at these redshifts.
In addition to providing spectroscopic confirmation of galaxies at the redshift frontier, deep prism spectra of distant galaxies have already illustrated the capability of providing tight constraints on stellar population parameters, finding very young ages (a few tens of Myr), very low metallicities (down to ), high yields of ionizing photons, low dust attenuation (Curtis-Lake et al., 2023; Atek et al., 2023). Yet, some of these constraints are still highly uncertain and often limited to relatively massive galaxies. These ultra-deep spectra will provide much tighter constraints on these parameters and push to much lower masses. For about 50 galaxies at z5 with S/N10 per resolution element on the continuum, it will even be possible to derive star formations histories (SFHs Looser et al., 2023a, b), hence testing the highly stochastic behavior of primeval galaxies expected by some models (Ceverino et al., 2018).
JWST spectroscopy has opened the exploration of the realm of intermediate mass black holes at high redshift, by detecting tens of AGN hosting black holes in the mass range between and , well below the BH mass range probed by quasar surveys (Furtak et al., 2023; Greene et al., 2023; Kocevski et al., 2023; Kokorev et al., 2023; Maiolino et al., 2023a; Matthee et al., 2023b; Übler et al., 2023). These studies have detected BHs in AGN that are still relatively luminous ( erg/s), as they were limited by the capability of detecting the nebular lines that can distinguish AGN from star forming galaxies. The ultra-deep spectra in program 3215 will allow the detection of even weaker active nuclei in early galaxies. Such AGNs can be recognized by their high ionization emission lines (NeV, NeIV, HeII), as well as other strong line diagnostics (Brinchmann, 2023; Nakajima & Maiolino, 2022; Tozzi et al., 2023). Further, in unobscured cases, we expect to detect the broad component of the H and H lines in grating spectra, separating this from gas outflows by comparing to the strong [OIII] lines. Comparing the inferred black hole masses to the host galaxy stellar mass will shed light on the formation channel of the first generation of black holes, differentiating between scenarios such as direct collapse, merging nuclear clusters, and Population III remnants (Habouzit et al., 2022; Volonteri et al., 2022; Pacucci et al., 2023; Schneider et al., 2023; Trinca et al., 2023).
@@
Target selection for the program 3215 spectroscopy will be described fully in a later paper. In brief, we have prioritized high-redshift candidates from JADES, progressively working to lower redshift. We include rare special targets at high priority, incorporating AGN indicators, quiescent galaxies at high redshift, sub-mm detections, and strong line emitters extracted from slitless spectroscopy from the FRESCO survey (Oesch et al., 2023). As reported in Table 2, we could allocate shutters to 228 unique targets. These are slightly fewer targets than allocated in program 1210, because in program 3215 we have given priority in re-observing the same targets in the five pointings, so to maximize the total exposure time for each target.
@@
Figure 4 shows the distribution of redshift and F444W magnitude for the targets that have been allocated shutters in the NIRSpec observation, illustrating that a significant number of them are very faint (). The color coding gives the total exposure time on each target (for the prism configuration), showing that many targets are observed for the full integration time of 168 ksec by being re-allocated shutters in the five pointings. We note that the three highest redshift galaxies from Curtis-Lake et al. (2023) are targeted for re-observation.
3.5 Additional Legacy Science
In addition to the primary science of probing the uncharted regime and revealing galaxy physics with ultra-deep spectroscopy in the parallel observations, the JOF data will provide a rich legacy data set of medium-band imaging and ultra-deep spectroscopy. We highlight here some examples of these opportunities.
The deep 1.6–4.1 m medium-band imaging will allow robust identification of quiescent galaxies via their Balmer breaks down to stellar mass. These data will constrain the evolution of the quiescent mass function and measure morphological gradients in Balmer break strengths (Ji et al., 2023), providing crucial constraints on the quenching mechanisms (e.g., Carnall et al., 2023; Nanayakkara et al., 2022; Santini et al., 2022; Marchesini et al., 2023; Looser et al., 2023a, b). The exquisite depth will also probe dwarf neighbors of larger mid-redshift galaxies, permitting analysis of their stellar populations and star-formation rates (Suess et al., 2023).
Turning to star-forming galaxies, by comparing the short-time-scale star-formation rates inferred from the measurement of H emission in the medium-bands to the longer time-scale values inferred from the rest-UV continuum, these data can test whether burstiness increases out to (e.g., Emami et al., 2019; Faisst et al., 2019; Reddy et al., 2023; Looser et al., 2023b). Bursts are hypothesized to be a critical driver of the strong emission lines observed at high- (e.g., Stark et al., 2013; Faucher-Giguère, 2018). With this very deep imaging, one can also map the line emission and hence star-formation rates across the face of these galaxies.
The data will also be an opportunity for the infrared time-domain, as the return to the deep Cycle 1 field with very deep year 2 imaging will reveal variability at micron on a timescale of year with medium-band resolution. This will probe supernovae (SNe) at (and potentially beyond), for which the SN emission shifts to micron with a peak brightness down to AB mag, as well as detect some AGN by their photometric variability.
@@ 3.6 Implementation
Program 3215 is implemented in the same manner as 1210, described in Eisenstein et al. (2023). NIRSpec must be the prime instrument with NIRCam in coordinated parallel, but by using multiple MSA configurations at slightly different pointings, we can achieve a 15-point dither pattern for NIRCam.
@@
We split the program into 5 visits, each with a pair of co-pointed MSA designs, one for the prism and one for the grating. Each visit contains 9 disperser/filter pairs, and each of these is observed for a 3-shutter nod. Each exposure contains 2 integrations for NIRSpec (19 groups of NRSIRS2 readout, Rauscher et al. 2017, 1400 s each) as well as 2 integrations for NIRCam (7 groups of DEEP8 readout, 1375 s each).
@@
In each of the 5 visits, we perform 4 nods with the G395M, 1 nod with G140M, and finally 4 nods with the PRISM. These are paired with 6 NIRCam exposures of F162M & F335M, 12 exposures of F182M & F250M, and 9 exposures of F210M & F300M. The total NIRCam exposure times are summarized in Table 1.
@@
The exact pointings of the 5 MSA pairs are chosen to place a small number of candidates onto all of the designs, while providing pixel diversity for both NIRCam and NIRSpec. We keep the total spacing small, under 2, as this increases the number of targets that can be observed on legal shutters in most of the designs. We also use the same position angle as 1210 (V3 PA of 321) and minimize the displacement of the two programs, to ensure high overlap of 14 filters.
As of this writing, program 3215 is scheduled to be observed in late October 2023, essentially on the one-year anniversary of the Year 1 observations of programs 1210 and 1180 in this field. Because program 3215 has no proprietary period, these data sets will go public at nearly the same time.

### v2

@@ 3.2 Observational Challenges and a Method to Reach
There are important reasons why pushing for F200W () dropouts is substantially harder than F115W ( or F150W () dropouts: 1) Such galaxies are fainter and rarer. 2) The frequency of extreme emission-line galaxies increases with redshift, and JWST clearly has shown a substantial population of them at , where they can create spectral energy distributions (SEDs) that differ from those of purely stellar populations, boosting the rest-optical and enhancing the apparent Balmer drop. 3) The dichroic gap in NIRCam leaves a wider spacing between F200W and F277W that makes it harder to exclude a mildly reddened Balmer-drop galaxy at , particularly with [OIII] potentially enhancing F277W. 4) Many Cycle 1 programs include only F277W, F356W, F410M, and F444W.111Notable exceptions are JEMS (C. C. Williams et al., 2023) and CANUCS (S. Withers et al., 2023). However, the filters beyond 4 microns have reduced sensitivity due to increased zodiacal foreground, so faint galaxies may only be well-detected in the bluer two bands, which in turn can be filled with [OIII] and H. These higher-redshift break candidates are less compelling than lower redshift examples that have 3 or 4 sensitive bands redward of the proposed break.
@@
Whereas Cycle 1 programs pushed deep in F115W and/or F150W, in program 3215 we adopt a strategy optimized to identify candidates at . In particular, this search must 1) push very deep in the 2 micron region; 2) increase the spectral resolution to isolate the dropout; and 3) include additional filters in the 2.5–4 micron region to eliminate contamination from mid-redshift emission-line galaxies. We expect that this combination of high-SNR detections, well-sampled dropout features, and systematic control over contaminants can reach the high standard of proof that claims of galaxies will require. The new imaging achieves these requirements by combining the usual NIRCam wide-band filters with well-chosen medium-band filters. By observing very deep in F182M, F210M, and F250M, the photometry can isolate a sharp break within a narrower and hence more distinctive region. Further, using the combination of medium and wide-bands in the long-wave (LW) channels reveals strong emission lines that could supply false positive Balmer breaks.
@@
We illustrate this strategy by considering the candidate in the CEERS field (C. T. Donnan et al., 2023), scaled down by a factor of 30 in flux and assuming the noise levels of programs 1210 & 3215. While now proved to be at (P. Arrabal Haro et al., 2023b), this object was a plausible candidate, showing a drop in F200W by a factor of 4 relative to F277W, no detection in F150W, and a blue continuum in the LW filters. However, this photometry was also a match to models at with strong optical emission lines and reddening suppression in the UV (R. P. Naidu et al., 2022b). Figure 2 shows how these two hypotheses are starkly differentiated by the NIRCam medium bands: the Lyman break is much steeper than a dust-enhanced Balmer break, and the medium bands easily distinguish continuum from emission lines.
The fits shown in Figure 2 to the scaled Cycle 1 wide-band photometry use the Prospector galaxy spectrum modeling code (B. D. Johnson et al., 2021). The Prospector models use FSPS (C. Conroy et al., 2009) with the MIST isochrones (J. Choi et al., 2016), the MILES stellar spectral library (P. Sánchez-Blázquez et al., 2006), and nebular emission following N. Byler et al. (2017). The fitted model includes a five-bin flexible SFH with broad priors on the ratio of SFR in adjacent bins following J. Leja et al. (2019); S. Tacchella et al. (2022) and normalization of a M. Kriek & C. Conroy (2013) dust attenuation curve. For the nebular emission, we adopt uniform priors on the log of the nebular ionization parameter and log of the gas-phase metallicity, which is allowed to vary independently of the stellar metallicity. Attenuation by the IGM follows P. Madau (1995). The resulting model space can produce a wide range of galaxy SEDs, making it a compelling test of the ability to reject low-redshift interpretations. We fit within several relatively narrow redshift ranges and tune the number of live points used for nested sampling with dynesty (J. S. Speagle, 2020; S. Koposov et al., 2022) to help identify narrow local minima in redshift, which can otherwise be missed (e.g. R. P. Naidu et al., 2022b).
Shifting the true galaxy in redshift and repeating the fits, Figure 3 demonstrates that this suite of filters provides very effective recovery of candidates at between 15 and 20. Here we show that for these reference galaxies, a strategy with medium bands produces a much larger rejection of the low-redshift hypothesis. This plot uses a fixed normalization of F277W signal-to-noise ratio, but brighter or fainter galaxies would simply scale as the square of the flux. What is important is the ratio of the between the two strategies, as this is proportional to how one would have the scale the exposure times in the background-noise limited case. One sees that at , this ratio is about 4 times higher, indicating that the medium-band strategy is more performant.
We note that these galaxy models are based on known stellar populations and do not include the extreme Balmer-break phenomenon associated to black-hole stars (R. P. Naidu et al., 2025a; A. de Graaff et al., 2025), plausibly connected to Little Red Dots (J. Matthee et al., 2024). Stronger Balmer breaks of course are harder to distinguish from high-redshift candidates, although medium-band photometry will help to distinguish a very abrupt Lyman- break from the smoother decline observed redward of Balmer break in these AGN (R. P. Naidu et al., 2025a; A. de Graaff et al., 2025). Such objects are rare and we don’t yet know about their abundance at faint flux levels. Importantly, they have strong rest-optical lines that should be readily measurable in spectra.
@@ 3.3 Filters, Exposure Times, and Depths
The driving requirement is to go deep in F182M and F210M, as potential dropout bands. In total, F182M, F200W, and F210M reached 91 hrs of exposure time, far more than the 11 hrs that 1210 & 1180 provide and much deeper than any other Cycle 1 program (to our knowledge, program 2079 being deepest with 17 hrs in F200W). We split the time roughly evenly between these two medium filters, as we do not know at what redshift a candidate might appear.
@@
By placing these observations on the parallel field created by program 1210, these new data build on the 55 hours of imaging depth already invested here. Given the steep number counts of galaxies at the redshift frontier, it is optimal to maximize the detection rate by integrating deeply on a single NIRCam pointing. The region in and around the HUDF in GOODS-S has the deepest and most comprehensive imaging and spectroscopy on the sky. While galaxies could be found anywhere, we site this program in GOODS-S to utilize and build upon the amazing legacy value of this field, while utilizing a position angle that allows ultra-deep MSA spectroscopy on the HUDF at a time of low zodiacal background. Of course, these medium-band images also support a wide range of other science cases for intermediate-redshift galaxies, notably emission-line morphologies and SED modeling around the Balmer break, as well as time-domain opportunities from the one-year cadence of the observations.
The results from this high-redshift search are summarized in §3.7, but we continue next with other aspects of the implementation of program 3215.
@@ 3.4 Coordinated Parallels with NIRSpec
| GOODS-S Year 2 | 3215 | 1 | 5 | 228 | 134 | 34 | 160 | ||
Program 3215 follows the path of 1210, using the NIRSpec MSA as the prime instrument and NIRCam as the coordinated parallel. This results in exceptionally deep spectroscopy on the HUDF and JADES Deep mosaic, about 91 hrs of total exposure, following on the 55 hrs of program 1210 spectroscopy released in A. J. Bunker et al. (2024).
Following the program goal of exploring the high-redshift universe, we optimize this time toward the study of high-redshift galaxies. Unlike program 1210, which was largely targeted from HST ACS optical and WFC3 infrared imaging (A. J. Bunker et al., 2024) save for four early high-redshift candidates from JADES (B. E. Robertson et al., 2023; E. Curtis-Lake et al., 2023), program 3215 utilizes NIRCam-selected targets. This allows a great expansion in the number of targets, as the JADES data is comfortably deeper than even the deepest WFC3 HUDF data and much deeper than the flanking CANDELS imaging.
Bringing ultra-deep spectroscopy to these faint high-redshift galaxies addresses many windows to understanding the physical processes associated with the early galaxy formation. We split the spectroscopic time across three dispersing modes—up to 134 ks on source with G395M (far deeper than program 1210), 160 ks with the Prism, and 34 ks with G140M—to serve a variety of key science drivers in studies of these galaxies.
@@
Finally, we seek to characterize the role of neutral hydrogen in and around the galaxies in the reionization epoch (e.g. G. C. Jones et al., 2024; A. Saxena et al., 2023; J. Witstok et al., 2024) by measuring the Ly emission line. The PRISM resolution is particularly coarse at the blue end of the spectrum, such that the grating substantially outperforms it for the detection of narrow lines. Further, because of the low resolution, the prism does not distinguish well between mild Ly emission and the Ly damping wing. The medium-deep G140M/F070LP spectra are therefore required and adequate to measure the Ly narrow line better than the prism and protect the damping wing analyses.
@@
The reductions of these ultradeep spectra were included in the JADES Data Release 3 (F. D’Eugenio et al., 2025a), along with details of the targeting. In brief, we have prioritized high-redshift candidates from JADES, progressively working to lower redshift. We included rare special targets at high priority, incorporating AGN indicators, quiescent galaxies at high redshift, sub-mm detections, and strong line emitters extracted from slitless spectroscopy from the FRESCO survey (P. A. Oesch et al., 2023). As reported in Table 2, we could allocate shutters to 228 unique targets. These are slightly fewer targets than allocated in program 1210, because in program 3215 we have given priority in re-observing the same targets in the five pointings, so to maximize the total exposure time for each target.
@@
Figure 4 shows the distribution of redshift and F444W magnitude for the targets that were allocated shutters in the NIRSpec observation, illustrating that a significant number of them are very faint (). The color coding gives the total exposure time on each target (for the prism configuration), showing that many targets are observed for the full integration time by being re-allocated shutters in the five pointings. We note that the three highest redshift galaxies from E. Curtis-Lake et al. (2023) are targeted for re-observation.
@@ 3.5 Implementation
Program 3215 was implemented in the same manner as 1210, described in D. J. Eisenstein et al. (2023). NIRSpec must be the prime instrument with NIRCam in coordinated parallel, but by using multiple MSA configurations at slightly different pointings, we can achieve a 15-point dither pattern for NIRCam.
@@
We split the program into 5 visits, each with a pair of co-pointed MSA designs, one for the prism and one for the grating. Each visit contains 9 disperser/filter pairs, and each of these is observed for a 3-shutter nod. Each exposure contains 2 integrations for NIRSpec (19 groups of NRSIRS2 readout, B. J. Rauscher et al. 2017, 1400 s each) as well as 2 integrations for NIRCam (7 groups of DEEP8 readout, 1375 s each).
@@
In each of the 5 visits, we perform 4 nods with the G395M, 1 nod with G140M, and finally 4 nods with the PRISM. These are paired with 6 NIRCam exposures of F162M & F335M, 12 exposures of F182M & F250M, and 9 exposures of F210M & F300M.
@@
The exact pointings of the 5 MSA pairs were chosen to place a small number of candidates onto all of the designs, while providing pixel diversity for both NIRCam and NIRSpec. We keep the total spacing small, under 2′′, as this increases the number of targets that can be observed on legal shutters in most of the designs. We also use the same position angle as 1210 (V3 PA of 321∘) and minimize the displacement of the two programs, to ensure high overlap of 14 filters.
3.6 Operational challenges
The 3215 program was observed October 16–24, 2023, with the data immediately released. Most of this data is of the hoped-for quality and has been used in numerous papers (e.g., B. Robertson et al., 2024; K. N. Hainline et al., 2024b; F. D’Eugenio et al., 2024; M. Curti et al., 2025). A minor data quality issue is that unsurprisingly we see a return of the blobs and oil slick artifacts that were seen in program 1210, described in (D. J. Eisenstein et al., 2023). We mask these in our reductions. We note that these features also appear (with small translations) in other GOODS-S pointings at this PA, even when the telescope is moved by over an arcminute, but in no pointings we have reduced at other PAs.
A recurring challenge with very deep NIRCam data is the print-through of background residuals from imperfect flat fields. We encountered this for the wide filters and reported in M. J. Rieke et al. (2023) that improved sky flats were needed to control the problem. Unsurprisingly, this also affects the medium-band images, and we have found that pushing to the depth of the JOF requires even better flats. We note that the dither pattern dictated by the NIRSpec, with 3-step nods and only small transverse steps between visits, appears to exacerbate the meso-scale residuals in the mosaics, even with 15 distinct pointings. We will discuss this in a future data release paper, showing that the issue is largely resolved at the depth of the JOF.
A more serious problem was that the 5th day of observations were impacted by short circuits in the NIRSpec MSA (T. D. Rawle et al., 2022). These badly flood the first 18 exposures (6 nods) of this visit, making useless all of the grating exposures and 3 of the 12 PRISM exposures. The remaining 9 PRISM exposures were not affected. The total NIRCam exposure times summarized in Table 1 do not include these impacted exposures.
This problem also created substantial stray-light banding in long-wave module A of the first 6 NIRCam exposures, all F335M. The other 12 short-impacted exposures do not show effects, perhaps because the filter changed to F250M, nor is there visible impact in the short-wave data. This banding must be masked.
The imaging observations also encountered substantial persistence in visits 1, 2, and 4. In visit 1, the first six exposures, all F162M, have diffuse persistence in modules A3, B3, and B4, typical of a large change in background. In visits 2 and 4, the A3 data are affected by persistence from prior wavefront sensing, revealing the pupil image and in visit 4 the ingress and egress trails of the target star. These effects are similar to those encountered in JADES (D. J. Eisenstein et al., 2023). As all of the artifacts described above affect only portions of the field, we leave the total exposure times unchanged in Table 1, but note that the effective exposure time on a given location can be mildly shorter.
We requested and were approved for reobservations of these 6 exposures, both NIRSpec and F162M+F335M imaging, which were planned to occur in October 2024 at the same PA. Unfortunately, this plan was rendered impossible due to an ill-timed telescope safing. Facing a delay to October 2025, we developed an alternative plan to instead use the remaining time to conduct spectroscopy on the JOF imaging and to use the NIRCam parallel to create a second shallower field with the same filter set. This will be described in § 3.8.
3.7 High-redshift results from the Origins Field
With the 2023 data, B. Robertson et al. (2024) analyzed the UV luminosity function using 9 galaxies in the JOF. Unfortunately, no galaxies were found, but the upper limits were used to contribute to the inference of the evolution of the UV luminosity function, measuring a drop by a factor of between and 14. N. J. Adams et al. (2025) performed a similar analysis in the field of both candidate recovery and the UV luminosity function.
However, the data set did reveal the astonishing JADES-GS-z14-0 (S. Carniani et al., 2024), now spectroscopically confirmed at (S. Carniani et al., 2025; S. Schouws et al., 2025a, b). This galaxy is very bright (45 nJy) and spatially extended, showing that it is not AGN-dominated. The JOF would have detected it easily a factor of 10 fainter or out beyond (B. Robertson et al., 2024). Importantly, the medium-band imaging in the JOF was central to selecting the object as a high-redshift candidate; its proximity to a nearby galaxy was confusing until the medium-band data clearly showed a steep dropout at a mildly different wavelength than the Balmer break of the neighbor. Selection was also encouraged by the deep JADES F770W MIRI imaging, which showed a photometric excess reasonably interpreted as the H-[OIII]5007Å emission-line complex (J. M. Helton et al., 2025).
With JADES-GS-z14-0 and the recent MoM-z14 (R. P. Naidu et al., 2025b), only twice fainter, it is clear that there is a detectable population of galaxies at . Indeed, the brightness of these two known galaxies might recommend a wider and shallower strategy. But whether bright or faint, the effects of the dichoic wavelength and the presence of strong emission lines in the LW filters, discussed in § 3.2, should be folded into the design and will likely continue to favor the use of medium bands.
The 3215 imaging data reduction and catalogs will be included in the upcoming JADES data release.
3.8 Dense-shutter spectroscopy on the JOF
With the opportunity to replan the remaining time into a window in December 2024, we proposed and were approved to change the pointing so as to conduct multi-object spectroscopy on the JOF. At the same time, the NIRCam parallels at this position angle (V3_PA 39∘) fell on deep wide-band imaging from JADES program 1286. By doing the same medium-band filter set, we could create a second field with the same 15 filters as the JOF, albeit not as deep. This was conducted on December 17–18, 2024.
This final pointing consists of a single MSA configuration, which was then observed for 9 nods of a 3-high shutter, providing 3 distinct dither locations. 8 of the nods use long exposures, the same as the original 3215 plan; the 9th is much shorter, only a single integration per position with 6 groups of NIRCam MEDIUM8 readout and 10 groups of NIRSpec NRSIRS2 readout, in order to use the available time allocation. The NIRCam imaging provides 3 long nods of F250M, F300M, and F210M, 2 long nods of F162M, 2 long and 1 short nod of F182M, and 1 long and 1 short nod of F335M, this last filter being kept short because JADES also provides some of this filter. We also began the visit with 1 long nod of F150W and F356W, arguing to duplicate this filter from program 1286 as a guard against time-variable sources, which we have found to present false positives for dropout galaxy searches in multi-epoch surveys (C. DeCoursey et al., 2025).
For the spectroscopy, we decided to attempt a more experimental use of the NIRSpec MSA. JADES had been routinely using designs that create overlapping grating spectra and non-overlapping prism spectra, relying on the prism spectra to reliably sort out which emission lines in the grating data belong to which target. Here, we opted to omit the prism spectra and observe a single MSA configuration with many overlapping grating spectra, relying on astrometric variations along the slit and priors from the photometric redshifts to associate the lines. In doing so, we achieve very high multiplex, 850 targets on this one configuration, with a focus solely on emission-line science. We open shutters on any target that is sufficiently well centered. The targets were galaxies with photometric redshifts above 3, whose strong emission lines would fall in the G235M and G395M grating, with flux limits down to 7 in the very deep LW imaging. We omit any galaxy brighter than 300 nJy, as those dispersed continua would be bright enough to add noise to the overlapping extractions.
The approach is motivated by several important technical facts. First, the NIRSpec grating spectra are detector-noise limited; it requires over 10 overlapping spectra to generate sufficient background to overcome this. This implies that overlapping the spectra does not generate a harmful amount of photon noise. Second, the astrometric knowledge from NIRCam gives a strong prior on where lines can occur along the 1.6′′ length of a 3-shutter slit. In the wavelength direction, one has additional ways to determine associations from detection of multiple lines and from priors from photometric redshifts and even the detection of the line in the multi-band imaging. This association problem is familiar from slitless spectroscopy, such as with the NIRCam grism; indeed, it is easier here because one know which targets were given a shutter. Third, while the density of faint galaxies is high enough that continuum spectra are overlapped, the density of detectable emission lines is low enough that one is very far from the confusion limit.
By using G235M and G395M together on one MSA configuration, we acquire 1.7–5.3 micron spectra, with only small wavelength gaps due to the chip gap. We observed 4 long nods in each of G235M and G395M, and a short nod in G395M, totaling 33.6 and 35.8 ksec, respectively. The spectral performance was very good, detecting lines down to 2– ergs/cm2/s, and we netted 539 successful redshifts in this one MSA configuration. Further description of the method, along with target selection, resulting performance, and the redshift catalog will be presented in D’Eugenio et al. (in prep.).